MODELING COMPETITIVE INTERACTIONS IN PROTEINS: VIBRATIONAL SPECTROSCOPY OF Na^{+}(N − MET HY LACET AMIDE)_{1}(H_{2}O)_{0−4} IN THE 3μm REGION

\underline{DOROTHY J MILLER}, NATALIE REBACZ, JAMES M LISY, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

To properly understand the preferred structures and biological properties of proteins, it is important to know how they interact with their aqueous environment. Competitive intra-peptide, peptide-water, ion-water, and ion-peptide interactions, including hydrogen bonding, play a key role in determining the structures and properties of proteins. The primary types of hydrogen bonding that occur are the intramolecular amide-amide (N − H ⋯ O = C′) and the intermolecular amide-water (O − H ⋯ O = C′ and H − O ⋯ H − N). N-methylacetamide (NMA), which contains both a carbonyl group and an amide group, is a common model for investigating these competitive interactions. An analysis of the infrared photo dissociation spectra of Na^{+}(NMA)_{1}(H_{2}O)_{0−4} in the O-H, N-H, and C=O spectral regions is presented. Parallel ab initio calculations are used as a guide in identifying both the type and location of non-covalent interactions present. In larger clusters, where several structural isomers may be present in the molecular beam, ab initio calculations are also used to suggest assignments for the observed IR transitions. The results presented offer insight to the nature of ion-NMA interactions present in an aqueous environment.